# Finite Sample Analysis of Distributional TD Learning with Linear Function Approximation

### v1

@@ Appendix F Analysis of the Categorical Projected Bellman Matrix
$r/\iota_{K}+\gamma j-K\leq 1/\iota_{K}+\gamma j-K\leq(1-\gamma)K+\gamma K-K=0<1.$ (39)
@@
Corollary F.1.
$\sum_{k=0}^{i}g_{j,k}(r)=\begin{cases}0,&\text{ for}\ 0\leq i<k_{j}(r),\\ 1-a_{j}(r),&\text{ for}\ i=k_{j}(r),\\ 1,&\text{ for}\ k_{j}(r)<i\leq K.\\ \end{cases}$ (40)
Corollary F.2.
For any $\gamma<1$, it holds that
$k_{j+1}(r)=\begin{cases}k_{j}(r),&\text{ if}\ a_{j}(r)\leq 1-\gamma,\\ k_{j}(r)+1,&\text{ if}\ a_{j}(r)>1-\gamma.\\ \end{cases}$ (41)
@@
$a_{j+1}(r)=\begin{cases}a_{j}(r)+\gamma,&\text{ if}\ a_{j}(r)\leq 1-\gamma,\\ a_{j}(r)+\gamma-1,&\text{ if}\ a_{j}(r)>1-\gamma.\\ \end{cases}$ (42)
Lemma F.2.
@@
$(\tilde{{\bm{G}}}(r))_{j,i}=g_{j,i}(r)-g_{K,i}(r)\quad\text{for }j,i\in\{0,1,\cdots,K-1\}.$ (43)
Using the previous corollaries, direct calculation show that if $k_{j+1}(r)=k_{j}(r)$,
$({\bm{C}}\tilde{{\bm{G}}}(r){\bm{C}}^{-1})_{j,i}=\sum_{k=0}^{i}g_{j,k}(r)-\sum_{k=0}^{i}g_{j+1,k}(r)=\begin{cases}0,&\text{ for}\ 0\leq i<k_{j}(r),\\ a_{j+1}(r)-a_{j}(r),&\text{ for}\ i=k_{j}(r),\\ 0,&\text{ for}\ k_{j}(r)<i<K.\\ \end{cases}$ (44)
@@
$({\bm{C}}\tilde{{\bm{G}}}(r){\bm{C}}^{-1})_{j,i}=\sum_{k=0}^{i}g_{j,k}(r)-\sum_{k=0}^{i}g_{j+1,k}(r)=\begin{cases}0,&\text{ for}\ 0<i<k_{j}(r),\\ 1-a_{j}(r),&\text{ for}\ i=k_{j}(r),\\ a_{j+1}(r),&\text{ for}\ i=k_{j+1}(r),\\ 0,&\text{ for}\ k_{j}(r)<i<K.\\ \end{cases}$ (45)
As a result, all entries in ${\bm{C}}\tilde{{\bm{G}}}(r){\bm{C}}^{-1}$ is non-negative. Moreover, the sum of each column and $\left\|{{\bm{C}}\tilde{{\bm{G}}}(r){\bm{C}}^{-1}}\right\|_{\infty}$ is $\gamma$ because
$\sum_{i=0}^{K-1}({\bm{C}}\tilde{{\bm{G}}}(r){\bm{C}}^{-1})_{j,i}=\begin{cases}a_{j+1}(r)-a_{j}(r)=\gamma,&\text{ if}\ k_{j+1}(r)=k_{j}(r),\\ 1-a_{j}(r)+a_{j+1}(r)=\gamma,&\text{ if}\ k_{j+1}(r)=k_{j}(r)+1.\\ \end{cases}$ (46)
The row sum of ${\bm{C}}\tilde{{\bm{G}}}(r){\bm{C}}^{-1}$ is
$\sum_{j=0}^{K-1}({\bm{C}}\tilde{{\bm{G}}}(r){\bm{C}}^{-1})_{j,i}=\sum_{m=0}^{i}g_{0,m}(r)-\sum_{m=0}^{i}g_{K,m}(r)\leq 1-0=1.$ (47)

### v2

@@ Appendix H Analysis of the Categorical Projected Bellman Matrix
$r/\iota_{K}+\gamma j-K\leq 1/\iota_{K}+\gamma j-K\leq(1-\gamma)K+\gamma K-K=0<1.$
@@
Corollary H.1.
$\sum_{k=0}^{i}g_{j,k}(r)=\begin{cases}0,&\text{ for}\ 0\leq i<k_{j}(r),\\ 1-a_{j}(r),&\text{ for}\ i=k_{j}(r),\\ 1,&\text{ for}\ k_{j}(r)<i\leq K.\\ \end{cases}$
Corollary H.2.
$k_{j+1}(r)=\begin{cases}k_{j}(r),&\text{ if}\ a_{j}(r)\leq 1-\gamma,\\ k_{j}(r)+1,&\text{ if}\ a_{j}(r)>1-\gamma.\\ \end{cases}$
@@
$a_{j+1}(r)=\begin{cases}a_{j}(r)+\gamma,&\text{ if}\ a_{j}(r)\leq 1-\gamma,\\ a_{j}(r)+\gamma-1,&\text{ if}\ a_{j}(r)>1-\gamma.\\ \end{cases}$
Lemma H.2.
@@
$(\tilde{{\bm{G}}}(r))_{j,i}=g_{j,i}(r)-g_{K,i}(r)\quad\text{for }j,i\in\{0,1,\cdots,K-1\}.$
Using the previous corollaries, through direct calculation we have that if $k_{j+1}(r)=k_{j}(r)$,
$({\bm{C}}\tilde{{\bm{G}}}(r){\bm{C}}^{-1})_{j,i}=\sum_{k=0}^{i}g_{j,k}(r)-\sum_{k=0}^{i}g_{j+1,k}(r)=\begin{cases}0,&\text{ for}\ 0\leq i<k_{j}(r),\\ a_{j+1}(r)-a_{j}(r),&\text{ for}\ i=k_{j}(r),\\ 0,&\text{ for}\ k_{j}(r)<i<K.\\ \end{cases}$
@@
$({\bm{C}}\tilde{{\bm{G}}}(r){\bm{C}}^{-1})_{j,i}=\sum_{k=0}^{i}g_{j,k}(r)-\sum_{k=0}^{i}g_{j+1,k}(r)=\begin{cases}0,&\text{ for}\ 0<i<k_{j}(r),\\ 1-a_{j}(r),&\text{ for}\ i=k_{j}(r),\\ a_{j+1}(r),&\text{ for}\ i=k_{j+1}(r),\\ 0,&\text{ for}\ k_{j}(r)<i<K.\\ \end{cases}$
As a result, all entries in ${\bm{C}}\tilde{{\bm{G}}}(r){\bm{C}}^{-1}$ is non-negative. Moreover, the sum of each column and $\left\|{{\bm{C}}\tilde{{\bm{G}}}(r){\bm{C}}^{-1}}\right\|_{\infty}$ is $\gamma$ since
$\sum_{i=0}^{K-1}({\bm{C}}\tilde{{\bm{G}}}(r){\bm{C}}^{-1})_{j,i}=\begin{cases}a_{j+1}(r)-a_{j}(r)=\gamma,&\text{ if}\ k_{j+1}(r)=k_{j}(r),\\ 1-a_{j}(r)+a_{j+1}(r)=\gamma,&\text{ if}\ k_{j+1}(r)=k_{j}(r)+1.\\ \end{cases}$
Moreover, the row sum of ${\bm{C}}\tilde{{\bm{G}}}(r){\bm{C}}^{-1}$ is
$\sum_{j=0}^{K-1}({\bm{C}}\tilde{{\bm{G}}}(r){\bm{C}}^{-1})_{j,i}=\sum_{m=0}^{i}g_{0,m}(r)-\sum_{m=0}^{i}g_{K,m}(r)\leq 1-0=1.$